OpenStudy (anonymous):

find the missing length. round to the nearest tenth if necessary x, 14, 9 A. 4.8 B. 5 C. 10.7 D. 11

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

use the pythagoren theorem

OpenStudy (anonymous):

\[a^{2}+b ^{2}=c^{2}\]

OpenStudy (anonymous):

a=9 c=14

OpenStudy (shamim):

14^2=9^2+x^2 x=?

OpenStudy (wolf1728):

x^2 = 14^2 -9^2

OpenStudy (mathmate):

@gamingnerd12 If you understand how the Pythagorian theorem works, you will be able to solve this problem and problem and probably thousands of problem related to it. Here's a definition and an article that will help you. In mathematics, the Pythagorean theorem, also known as Pythagoras's theorem, is a relation in Euclidean geometry among the three sides of a right triangle. It states that the square of the hypotenuse (the side opposite the right angle) is equal to the sum of the squares of the other two sides. http://www.mathsisfun.com/pythagoras.html Many problems later on in your math career will depend on it, so it's avantageous to spend an hour or two now to save tens of hours in the future.